OpenStudy (anonymous):

(1 - 5/x)^x with lim ->infinity

OpenStudy (anonymous):

is it e^5?

OpenStudy (amistre64):

as x gets large; 5/x goes to zero; which leaves 1^inf .... is say it goes to 1

OpenStudy (amistre64):

well, wolfram says it e^(-5)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=lim {x+to+infinity}%281+-+5%2Fx%29^x

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\(\infty^0\) is one of the standard indeterminate forms. For this you must employ l'Hopital's rule. http://en.wikipedia.org/wiki/Indeterminate_form#List_of_indeterminate_forms

OpenStudy (zarkon):

\[\lim_{x\to \infty}\left(1+\frac{r}{x}\right)^{xt}=e^{rt}\] replace r by -5 and t by 1

OpenStudy (anonymous):

Yeah, if you know all the different forms of e you can just do it by sight ;)